OpenStudy (anonymous):

How do you simplify a rational expression? For example.. 3x-6/x-2.

OpenStudy (ranga):

Factor out 3 from the numerator and simplify.

OpenStudy (anonymous):

I forgot how to factor...

OpenStudy (anonymous):

How do i factor out 3?

OpenStudy (ranga):

(3x - 6) /(x - 2) = 3(x - 2)/(x - 2) = 3